As of today, assume the following information is available:

                                                                        U.S.             Mexico

      Real rate of interest required by investors     1%                1%

      Nominal interest rate                                    2%                 6%

      Spot Rate (St)                          0.0560 USD/MXN

      One‑year forward rate (Ft,1-yr)   0.0538 USD/MXN

  1.     What is the expected inflation in the US and Mexico using fisher effect. [Hint: you can get the expected inflation using Fisher’s formula –i.e., the nominal interest rate is equal to the real interest rate plus expected inflation]  

  1.    Calculate the change in the forward and spot rate for USD/MXN. Compare that with the difference between the nominal interest rates in US and Mexico. Does forward expectation parity holds?

  1.     Suppose that, due to change in economic relation, the difference in expected inflation between US and Mexico  is -3.5% instead of 4%. What will be the revised nominal interest rate in Mexico using the fisher effect assuming the nominal interest rate in USA stay the same? [Hint: You can use the approximation method, i.e. difference in nominal interest rate is proportional to difference in expected inflation ]
Answer #1

(1+nominal rate)= (1+ real rate)*(1+inflation)

In US

(1+2%)=(1+1%)*(1+x)

1+x=1.02/1.01

x=(1.02/1,01)-1

x=1.009901-1

x=0.009901

There fore expected inflation rate is 0.99%

same  for Mexico

(1+6%)=(1+1%)*(1+x)

x=4.95%